OpenStudy (anonymous):

3. Given the following polynomial function: f(x) = 3x^4 + 2x^3 – 2 b. Find critical value(s) c. Find interval(s) where f(x) is increasing. d. Find interval(s) where f(x) is decreasing e. State any Max or min. f. Find interval(s) where f(x) is concave up g. Find interal(s) where f(x) is concave down h. State any point(s) of inflection

OpenStudy (anonymous):

\[y = 3x ^{4}+2x ^{3}-2\]

OpenStudy (anonymous):

For critical values take dy/dx =0

OpenStudy (anonymous):

\[\frac{dy}{dx}=12x^3+6x^2\]

OpenStudy (anonymous):

\[\frac{d^2y}{dx^2}=24x^2+12x\]

OpenStudy (anonymous):

how do i find the critical value

OpenStudy (anonymous):

after getting the derivative

OpenStudy (anonymous):

0 sorry about the answer earlier

OpenStudy (anonymous):

\[0=12x^3+6x^2\]

OpenStudy (anonymous):

it says state any points of inflection have any clue?

OpenStudy (anonymous):

do the same to the second derivative. its critical points are usually inflection points.

OpenStudy (anonymous):

when a first derivative is positive, it's original function is increasing. negative, decreasing when it is zero, the original equation is flat (like at a minimum or maximum.

OpenStudy (anonymous):

second derivative: if it is positive, the original function is concave up negative, concave down. zero = inflection point

OpenStudy (anonymous):

notice a difference between increasing and concave up. a curve can be decreasing, but doing so at a slower and slower rate, thus concave up.